the base of a triangular sail is 20m shorter than it's height.
the area is 150m2. find the base and height of the sail.

Answer by unlockmath(1688)

Hello,
Let's represent the height of the sail with x and the base will be x-20.
We know the area of a triangle is 1/2(Base)(height), so we can plug in the values given.
1/2(x)(x-20)=150m2. Let's multiply by 2 to remove the fraction.
(x)(x-20)=300m2 Now multiply to give us:
x^2-20x=300m2 Now subtract 300 from both sides.
x^2-20x-300=0 This can be factored into:
(x-30)(x+10)=0 this gives us
x=30 and x=-10
30 Makes sense so the height is 30 meters and the base is 10 meters.
RJ Toftness
www.math-unlock.com